# Temporal Fair Division of Indivisible Goods with Scheduling

### v1

@@ A.9 Proof of Theorem 10
Assume there exists a solution that computes a $k$-TEFX allocation. Consider the instance with two agents and three goods $O=\{g_{1},g_{2},g_{3}\}$, where agents have identical valuations: $v(g_{1})=v(g_{2})=1$ and $v(g_{3})=\frac{1}{k}$, where $k>0$ and is independent of the values of the goods. In order to obtain a $k$-TEFX allocation at the first round, $g_{1}$ and $g_{2}$ must be allocated to each agent, respectively. Then, no matter which agent is allocated $g_{3}$, we do not obtain a $k$-TEFX allocation. Without loss of generality, let it be agent $1$, we have
$v_{2}(g_{2})=1<k\cdot v_{1}(g_{1}\cup g_{3})=k\cdot(1+\frac{1}{k})=k+1\text{,}$
@@ A.10 Algorithm and Proof of Theorem 11
We use MAX-ECE, which returns a $\frac{1}{2}$-EFX allocation (Theorem 3). Intuitively, we consider the first $n$ goods as a single round and run MAX-ECE. This is an EFX allocation because each agent gets exactly one good, which must be valued positively for all agents, and the removal of this good eliminates any envy towards any agent. Then, for each remaining round, we compute using MAX-ECE, and the final allocation is simply the union of all previous rounds’ allocations. It is noted that we cannot use the Draft-and-Eliminate algorithm by [5], which returns a ($\phi-1$)-EFX allocation, as it cannot guarantee that each agent is first allocated exactly one good, and if we forcefully do so, it may break the approximation guarantee.
@@
A.11 Algorithm and Proof of Theorem 12
@@
A.12 Proof of Lemma 1
@@
First, since the original algorithm maintains TEF1, agent $j$ does not envy agent $i$ before the allocation of good $g$.
Case 1: Agent $j$’s allocation does not have any positively valued good for themselves. Assume agent $i$ has at least one zero-valued good to agent $j$. Let $g_{1}$ be a good such that $v_{i}(g_{1})=b$, $v_{j}(g_{1})=0$, and good $g$ is a good such that $v_{i}(g)=v_{j}(g)=b$. This is because agent $i$ cannot be agent lastlast (Thus, the zero-valued good for agent $j$ is not valued zero for agent $i$), and the algorithm must allocate the good to an agent that values it positively. It means
$v_{i}(A^{t}_{i})\geq v_{i}(g_{1})+v_{i}(g)>0=v_{j}(A^{t}_{j})\text{.}$
Thus, good $g$ should not be allocated to agent $i$. Therefore, agent $i$ does not have any zero-valued good to agent $j$.
Case 2: Agent $i$ has at least one zero-valued good to agent $j$. Assume agent $j$’s allocation does not have any positively valued good for themselves. If agent $i$ is agent lastlast, then it contradicts our assumption, where agent $j$ must have obtained at least one positively valued good for themselves. If agent $i$ is not agent lastlast, it means agent $i$ has at least two goods. Let $g_{1}$ be a good such that $v_{i}(g_{1})=b$, $v_{j}(g_{1})=0$, and good $g$ is a good such that $v_{i}(g)=v_{j}(g)=b$. Then we have returned to the proof for Case 1, where we obtain a contradiction. Therefore, agent $j$’s allocation has at least one positively valued good for themselves. ∎
A.13 Algorithm and Proof of Theorem 14
@@ A.14 Proof of Theorem 15
A.15 Proof of Theorem 16
@@
A.16 Proof of Lemma 2
@@
A.17 Proof of Lemma 3
@@
A.18 Proof of Theorem 18
@@
Consider the instance with two agents and three goods $O=\{g_{1},g_{2},g_{3}\}$, where agents have identical valuations: $v(g_{1})=v(g_{2})=1$ and $v(g_{3})=2$ [14], and $r\leq T-1$. $O_{1}=\{g_{1},g_{2}\}$, $\forall 1<t<T,O_{t}=\{\emptyset\}$, and $O_{T}=\{g_{3}\}$. Essentially, we insert rounds in which no goods arrive, such that no matter how the goods are scheduled, $g_{1}$ and $g_{2}$ must be allocated before $g_{3}$.
To prove that TEFX does not exist in this instance, we encourage interested readers to read the proof in [14], as our proof closely resembles it. Specifically, when $T=2$, $r\leq 1$, which is the case where there are no additional rounds, it is equivalent to the case in their proof. ∎
A.19 Algorithm and Proof of Theorem 19
Algorithm 13 Returns a TEFX allocation under Identical Days when $n=2$
Output: TEFX allocation $\mathcal{A}$
1:if $T$ mod $4=0$ or $T$ mod $4=1$ then
2:  for $t=1$ to $\lfloor\frac{T}{2}\rfloor-1$ do
3:   for $g\in O_{t}$ do
4:     $delay(g,t,\lfloor\frac{T}{2}\rfloor-t+1)$
5:   end for
6:  end for
7:  Allocate equal copies of each good to agent $1$ and $2$ in round $\lfloor\frac{T}{2}\rfloor$.
8:else
9:  for $t=1$ to $\lfloor\frac{T}{2}\rfloor-2$ do
10:   for $g\in O_{t}$ do
11:     $delay(g,t,(\lfloor\frac{T}{2}\rfloor-1)-t+1)$
12:   end for
@@ A.19 Algorithm and Proof of Theorem 19
14:  Allocate equal copies of each good to agent $1$ and $2$ in round $\lfloor\frac{T}{2}\rfloor-1$.
15:  $m\leftarrow|O_{\lfloor\frac{T}{2}\rfloor}|$
16:  for $l=1$ to $m$ do
17:   if $v_{1}(A^{\lfloor\frac{T}{2}\rfloor}_{1})<v_{1}(A^{\lfloor\frac{T}{2}\rfloor}_{2})$ and $v_{2}(A^{\lfloor\frac{T}{2}\rfloor}_{2})<v_{2}(A^{\lfloor\frac{T}{2}\rfloor}_{1})$ then
18:     Swap $A^{\lfloor\frac{T}{2}\rfloor}_{1}\backslash A^{{\lfloor\frac{T}{2}\rfloor}-1}_{1}$ and $A^{\lfloor\frac{T}{2}\rfloor}_{2}\backslash A^{{\lfloor\frac{T}{2}\rfloor}-1}_{2}$
19:   end if
20:   Let $i$ be an unenvied agent
21:   Let $g^{*}\in argmax_{g\in O_{\lfloor\frac{T}{2}\rfloor}}v_{1}(g)$
22:   $A^{\lfloor\frac{T}{2}\rfloor}_{1}\leftarrow A^{\lfloor\frac{T}{2}\rfloor}_{1}\cup\{g^{*}\}$
[…]
24:  end for
25:  if $v_{2}(A^{\lfloor\frac{T}{2}\rfloor}_{2})<v_{2}(A^{\lfloor\frac{T}{2}\rfloor}_{1})$ then
26:   Swap $A^{\lfloor\frac{T}{2}\rfloor}_{1}\backslash A^{{\lfloor\frac{T}{2}\rfloor}-1}_{1}$ and $A^{\lfloor\frac{T}{2}\rfloor}_{2}\backslash A^{{\lfloor\frac{T}{2}\rfloor}-1}_{2}$
27:  end if
28:end if
29:if $T$ mod $4=0$ or $T$ mod $4=2$ then
30:  for $t=\lfloor\frac{T}{2}\rfloor+1$ to T1T-1 do
31:   for $g\in O_{t}$ do
32:     delay(g,t,Tt+1)delay(g,t,T-t+1)
33:   end for
34:  end for
35:  Allocate equal copies of each good to agent $1$ and $2$ in round $T$.
36:else
37:  for $t=\lfloor\frac{T}{2}\rfloor+1$ to T2T-2 do
38:   for $g\in O_{t}$ do
39:     delay(g,t,(T1)t+1)delay(g,t,(T-1)-t+1)
40:   end for
41:  end for
42:  Allocate equal copies of each good to agent $1$ and $2$ in round T1T-1.
43:  $m\leftarrow|O_{T}|$
44:  for $l=1$ to $m$ do
45:   if $v_{1}(A^{T}_{1})<v_{1}(A^{T}_{2})$ and $v_{2}(A^{T}_{2})<v_{2}(A^{T}_{1})$ then
46:     Swap $A^{T}_{1}\backslash A^{T-1}_{1}$ and $A^{T}_{2}\backslash A^{T-1}_{2}$
47:   end if
48:   Let $i$ be an unenvied agent
49:   Let $g^{*}\in argmax_{g\in O_{T}}v_{1}(g)$
50:   $A^{T}_{1}\leftarrow A^{T}_{1}\cup\{g^{*}\}$
51:   $OT\{$$\leftarrow O_{T}\backslash\{g^{*}\}$
52:  end for
53:  if $v_{2}(A^{T}_{2})<v_{2}(A^{T}_{1})$ then
54:   Swap $A^{T}_{1}\backslash A^{T-1}_{1}$ and $A^{T}_{2}\backslash A^{T-1}_{2}$
55:  end if
56:end if
Proof.
We prove Theorem 19 by considering cases according to the divisibility of $T$ on $4$.
Case 1: $T$ mod $4=0$. In round $\frac{T}{2}$, we allocate the same number of copies of goods to both agents. Thus, we obtain a TEFX (TEF) allocation. Then, in round $T$, we also allocate the same number of copies of goods to both agents.
$v_{1}(A_{1})=v_{1}(A_{2}),v_{2}(A_{1})=v_{2}(A_{2})\text{.}$
Therefore, we obtain a TEFX (TEF) allocation.
Case 2: $T$ mod $4=2$. In round $\frac{T}{2}-1$, there is an even number of copies of each good, so we are able to allocate the same number of copies of goods to both agents, resulting in a TEFX (TEF) allocation. In round $\frac{T}{2}$, intuitively, we run MAX-ECE to compute an EFX allocation (Theorem 2). A key part of this is that instead of swapping the entire allocation, we only need to swap the goods that are allocated in this round. Denote the set of goods $P$ to be $A^{\frac{T}{2}}_{2}\backslash A^{\frac{T}{2}-1}_{2}$. Notice that
$v_{1}(A^{T}_{2})=v_{1}(P\cup A^{\frac{T}{2}-1}_{2})=v_{1}(P\cup A^{\frac{T}{2}-1}_{1})\text{.}$
A similar proof can be obtained for agent $2$. Thus, in round $\frac{T}{2}$, we obtain a TEFX allocation. Then, in round $T$, for each agent, we allocate the copies of good such that each agent has the same number of copies of goods, which results in a TEFX (TEF) allocation, since in round $\frac{T}{2}$, each agent either obtains $\frac{T}{4}-1$, or $\frac{T}{4}$ copies of good.
Case 3: $T$ mod $4=1$. $\lfloor\frac{T}{2}\rfloor$ is even. Thus, for round $1$ to T1T-1, it is equivalent to Case 1. Now, each agent has the same number of copies of goods. In round $T$, we run MAX-ECE again as in Case 2, resulting in a TEFX allocation.
Case 4: $T$ mod $4=3$. Note that $\lfloor\frac{T}{2}\rfloor$ is odd. Thus, for round $1$ to T1T-1, it is equivalent to Case 2. Now, each agent has the same number of copies of goods. In round $T$, we run MAX-ECE again as in Case 2, resulting in a TEFX allocation. ∎
A.20 Proof of Theorem 20
Proof.
We use the same instance from Theorem 18. For completeness, we state the instance here again: Consider the instance with two agents and three goods $O=\{g_{1},g_{2},g_{3}\}$, where agents have identical valuations: $v(g_{1})=v(g_{2})=1$ and $v(g_{3})=2$ [14], and $r\leq T-1$. $O_{1}=\{g_{1},g_{2}\}$, $\forall 1<t<T,O_{t}=\emptyset$, and $O_{T}=\{g_{3}\}$. As mentioned in the proof of Theorem 18, we insert rounds in which no goods arrive, such that no matter how the goods are scheduled, $g_{1}$ and $g_{2}$ must be allocated before $g_{3}$.
In round $1$, to obtain a TMMS allocation, each agent must be allocated either $g_{1}$ or $g_{2}$. Without loss of generality, assume agent 11 is allocated $g_{3}$. However, notice that in order to obtain a TMMS allocation, each agent must be allocated either $\{g_{1},g_{2}\}$ or $g_{3}$, which is impossible to achieve. ∎

### v3

@@ 4.3.1 General Settings
Assume there exists a solution that computes a $k$-TEFX allocation. Consider the instance with two agents and three goods $O=\{g_{1},g_{2},g_{3}\}$, where agents have identical valuations: $v(g_{1})=v(g_{2})=1$ and $v(g_{3})=\frac{1}{k}+1$, where $k>0$ and is a constant. Let $g_{i}$ arrive in round $i$. In order to obtain a $k$-TEFX allocation at the first round, $g_{1}$ and $g_{2}$ must be allocated to each agent, respectively. Then, no matter which agent is allocated $g_{3}$, we do not obtain a $k$-TEFX allocation. Without loss of generality, let it be agent $1$, we have
$v(g_{2})=1<k\cdot(v(g_{1}\cup g_{3})-v(g_{1}))=k\cdot(\frac{1}{k}+1)=k+1\text{,}$
@@
We use MAX-ECE, which returns a $\frac{1}{2}$-EFX allocation (Theorem 3). Intuitively, we consider the first $n$ goods as a single round and run MAX-ECE. This is an EFX allocation because each agent gets exactly one good, which must be valued positively for all agents, and the removal of this good eliminates any envy towards any agent. Then, for each remaining round, we compute using MAX-ECE, and the final allocation is simply the union of all previous rounds’ allocations. It is noted that we cannot use the Draft-and-Eliminate algorithm by Amanatidis et al. (2020), which returns a ($\phi-1$)-EFX allocation, as it cannot guarantee that each agent is first allocated exactly one good, and if we forcefully do so, it may break the approximation guarantee.
@@
4.3.2 Identical Days
Theorem 11.
@@ 4.3.2 Identical Days
4.3.3 Generalized Binary Valuation
Theorem 12.
@@ 4.3.3 Generalized Binary Valuation
Algorithm 6 is a modified version of Algorithm 3 from Elkind et al. (2025), which returns a TEF1 allocation under the generalized binary valuation setting.
We first show a key property for Algorithm 6.
Lemma 7.
Output: $\frac{1}{2}$-TEFX allocation $\mathcal{A}$
$1last\leftarrow$-1
2:for $t=1$ to $T$ do
3:  for $g\in O_{t}$ do
4:   $S\leftarrow\{i^{*}\in N|v_{i^{*}}(g)>0\}$
5:   if $S\neq\emptyset$ then
7:     if $A^{t}_{i}=\emptyset$ then
8:      $last\leftarrow i$
9:     end if
10:     $A^{t}_{i}\leftarrow A^{t}_{i}\cup\{g\}$
11:     $Ot\{$$\leftarrow O_{t}\backslash\{g\}$
12:   end if
13:  end for
14:end for
15:for $i\in N$ do
16:  if $A_{i}=\emptyset$ then
17:   $last\leftarrow i$
18:  end if
19:end for
20:for $t=1$ to $T$ do
21:  for $g\in O_{t}$ do
22:   $A^{t}_{last}\leftarrow A^{t}_{last}\cup\{g\}$
23:  end for
24:end for
4.3.4 Identical Valuation
Theorem 13.
@@ 4.3.5 Bi-valued goods
In summary, our exploration of $\alpha$-TEFX demonstrates that while exact temporal envy-freeness is overwhelmingly brittle in dynamic environments, multiplicative approximations offer a robust and mathematically viable alternative. We established that universal constant-factor guarantees are generally impossible due to inherent valuation variance and the specific challenges posed by zero-valued goods. However, by introducing structural constraints such as the periodicity of identical days or the rigid utility tiers of generalized binary valuations, we successfully identified scenarios in which strong constant-factor approximations, such as $1/2$-TEFX, can be recovered. Even in environments where constant bounds remain elusive, such as identical valuation settings, carefully designed greedy mechanisms allow us to mitigate extreme disparities and secure optimized, value-dependent guarantees. Ultimately, these results confirm that relaxing the fairness criteria from exact to multiplicative allows us to bridge the gap between theoretical impossibility and practical, algorithmically sound fair division.
4.4 TMMS
In this section, we consider Temporal Maximin Share Fairness (TMMS). Note that MMS does not exist when there are more than two agents (Kurokawa et al., 2016, 2018). We are the first to consider MMS under the temporal fair division model.
Here, we note that under the generalized binary valuation setting, EF1 implies MMS (Bouveret and Lemaître, 2016). Therefore, since Elkind et al. (2025) proved the possibility for TEF1 under the generalized binary valuation setting, TMMS also exists under this setting.
4.4.1 Identical Days
Theorem 15.
@@ 4.4.1 Identical Days
In summary, our investigation into Temporal Maximin Share Fairness (TMMS) reveals that the cumulative nature of maximin guarantees is exceptionally difficult to satisfy in a dynamic context. Unlike static settings, where MMS is often a more accessible benchmark than EFX, the requirement to meet the maximin threshold in every round makes TMMS remarkably fragile. We demonstrated that even with only two agents and identical valuations, the sequential arrival of goods, whether in general settings or repeating identical days, frequently precludes the existence of a TMMS allocation. The sole sanctuary of this fairness notion lies in the highly structured domain of generalized binary valuations, where the existence of TEF1 suffices for a positive result. Ultimately, these findings suggest that while TMMS is a theoretically interesting extension of the maximin principle, it is often too restrictive for general temporal allocation, reinforcing the need for the approximation-based or preference-restricted approaches we have explored.
5 Temporal Fair Division with Scheduling
5.1 TEF1
5.1.1 Identical Days
Theorem 16.
Algorithm 8 $R$’
Output: EF1 allocation $\mathcal{A}$
1:for $i\in N$ do
2:  $A_{i}\leftarrow\emptyset$
3:end for
4:for $l=1$ to $m$ do
5:  $i\leftarrow(i$ mod $m)+1$
7:  $A_{i}\leftarrow A_{i}\cup\{g\}$
8:  $M\{$$\leftarrow M\backslash\{g\}$
9:end for
Proof.
Denote $\lambda=n$.
Algorithm 9 Returns a TEF1 allocation when $r\geq\frac{n}{2}$
Output: TEF1 allocation $\mathcal{A}$
[…]
3:for $g\in O_{1}$ do
4:  Append $g$ to cpycpy
5:end for
[…]
7:while $k\cdot n+n-1<T$ do
8:  for $t=0$ to $\lceil\frac{n}{2}\rceil-1$ do
9:   for $g\in O_{t+k}$ do
10:     $delay(g,t+k,(k\cdot n+\lceil\frac{n}{2}\rceil)-(t+k)+1)$
11:   end for
12:  end for
[…]
14:  for $t=\lceil\frac{n}{2}\rceil+1$ to n1n-1 do
15:   for $g\in O_{t+k}$ do
16:     $delay(g,t+k,(k\cdot n+n-1)-(t+k)+1)$
17:   end for
18:  end for
19:  Allocate $O_{k\cdot n+n-1}$ such that
[…]
21:  $k\leftarrow k+1$
22:end while
23:for $t=k$ to $\frac{T+k}{2}-1$ do
24:  for $g\in O_{t}$ do
25:   $delay(g,t,\frac{T+k}{2}-t+1)$
26:  end for
27:end for
[…]
29:for $t=\frac{T+k}{2}+1$ to T1T-1 do
30:  for $g\in O_{t}$ do
31:   delay(g,t,Tt+1)delay(g,t,T-t+1)
32:  end for
33:end for
34:Reverse NN
[…]
Intuitively, Algorithm 9 schedules all of the goods such that every $\lambda$ rounds, it is scheduled into two phases, where in each phase it contains approximately $\frac{\lambda}{2}$ rounds’ cumulative goods. We attempt to maintain a state where, after every two phases ($\lambda$ rounds), each agent has exactly one copy of each good. To achieve this, for the first phase, we need to obtain an EF1 allocation, ensuring that no agent receives more than one copy of each good. At last, if $T$ is not divisible by $\lambda$, we schedule the goods into two rounds, applying the result from Elkind et al. (2025) to compute a TEF1 allocation. We note that the result from Elkind et al. (2025) cannot be applied to allocating scheduled phases, as our result relies on the fact that in every $\lambda$ rounds, we obtain a TEF allocation, whereas their result does not guarantee this.
We show a key property in Algorithm 8.
Lemma 10.
@@ 5.1.1 Identical Days
Next, we show a key property in Algorithm 9.
Lemma 11.
@@
[…]
Lastly, if $T$ is not divisible by $\lambda$, we can schedule the remaining rounds as two rounds as $r\geq\frac{\lambda}{2}\geq\frac{T}{2}$. By Elkind et al. (2025), by running an RR in the first scheduled round, and another RR with the allocation order reversed in the second (last) scheduled round, we obtain a TEF1 allocation. ∎
5.2 TEFX
Theorem 17.
@@ 5.2 TEFX
Consider the instance with two agents and three goods $O=\{g_{1},g_{2},g_{3}\}$, where agents have identical valuations: $v(g_{1})=v(g_{2})=1$ and $v(g_{3})=2$, and $r\leq T-1$. $O_{1}=\{g_{1},g_{2}\}$, $\forall 1<t<T,O_{t}=\{\emptyset\}$, and $O_{T}=\{g_{3}\}$. Essentially, we insert rounds in which no goods arrive, such that no matter how the goods are scheduled, $g_{1}$ and $g_{2}$ must be allocated before $g_{3}$.
To prove that TEFX does not exist in this instance, we encourage interested readers to read the proof in Elkind et al. (2025), as our proof closely resembles it. Specifically, when $T=2$, $r\leq 1$, which is the case where there are no additional rounds, it is equivalent to the case in their proof. ∎
5.3 TMMS
Theorem 18.
Proof.
We use the same instance from Theorem 17. For completeness, we state the instance here again: Consider the instance with two agents and three goods $O=\{g_{1},g_{2},g_{3}\}$, where agents have identical valuations: $v(g_{1})=v(g_{2})=1$ and $v(g_{3})=2$, and $r\leq T-1$. $O_{1}=\{g_{1},g_{2}\}$, $\forall 1<t<T,O_{t}=\emptyset$, and $O_{T}=\{g_{3}\}$. As mentioned in the proof of Theorem 17, we insert rounds in which no goods arrive, such that no matter how the goods are scheduled, $g_{1}$ and $g_{2}$ must be allocated before $g_{3}$.
In round $1$, to obtain a TMMS allocation, each agent must be allocated either $g_{1}$ or $g_{2}$. Without loss of generality, assume agent $1$ is allocated $g_{3}$. However, notice that in order to obtain a TMMS allocation, each agent must be allocated either $\{g_{1},g_{2}\}$ or $g_{3}$, which is impossible to achieve. ∎
6 Discussion
Several open questions remain regarding the extension of the temporal fair division framework. One key question is to tighten the scheduling bounds. Our result establishes $r\geq T/2$ as a buffer for TEF1, which may not be a tight constraint. It remains an open question whether a smaller scheduling buffer can be obtained.
Appendix A Appendix
A.1 Pseudocode of ECE
Algorithm 10 ECE
Output: Allocation $\mathcal{A}$
1:for $i\in N$ do
2:  $A_{i}\leftarrow\emptyset$
3:end for
4:for $l=1$ to $m$ do
5:  while there does not exist an unenvied agent do
6:   Find an envy-cycle $C=\left(i_{0},\dots,i_{d-1}\right)$
7:   $\forall 0\leq j<d$, assign the allocation of $i_{j+1\text{mod}d}$ to $i_{j}$.
8:  end while
9:  Let $i$ be an unenvied agent
10:  Let $g^{*}$ be an arbitrary good
11:  $A_{i}\leftarrow A_{i}\cup\{g^{*}\}$
12:  $M\{$$\leftarrow M\backslash\{g^{*}\}$
@@ A.1 Pseudocode of ECE
A.2 Psuedocode of MAX-ECE
Algorithm 11 MAX-ECE
Output: Allocation $\mathcal{A}$
1:for $i\in N$ do
2:  $A_{i}\leftarrow\emptyset$
3:end for
4:for $l=1$ to $m$ do
5:  while there does not exist an unenvied agent do
6:   Find an envy-cycle $C=\left(i_{0},\dots,i_{d-1}\right)$
7:   $\forall 0\leq j<d$, assign the allocation of $i_{j+1\text{mod}d}$ to $i_{j}$.
8:  end while
9:  Let $i$ be an unenvied agent
10:  Let $g^{*}=argmax_{g\in M}v_{i}(g)$
11:  $A_{i}\leftarrow A_{i}\cup\{g^{*}\}$
12:  $M\{$$\leftarrow M\backslash\{g^{*}\}$
13:end for
A.3 Psuedocode of RR
Algorithm 12 $R$
Output: Allocation $\mathcal{A}$
1:for $i\in N$ do
2:  $A_{i}\leftarrow\emptyset$
3:end for
4:for $l=1$ to $m$ do
5:  Let $i\leftarrow N_{l\text{mod}n}$
6:  Let $g^{*}\in argmax_{g\in M}v_{i}(g)$
7:  $A_{i}\leftarrow A_{i}\cup\{g^{*}\}$
8:  $M\{$$\leftarrow M\backslash\{g^{*}\}$
9:end for